How to solve questions in thermodynamic using 1 law of thermodynamics Give some tricks

Grade:11

1 Answers

naveen
6 years ago
Given DQ=du+dwDQ=2000heat energy supplyDw=pdv pressure and change in volumePdv=2×10Du=D-Day2000-(2×10×0.004)1200 joules